[5] 20 Find the partial derivatives home and my f(x,y) = (xy - 12
Answer #1

2 в у = (-1) OF ә/ (- ) 2 -(xy-1) (nyol) dr up dx df = 2 (2-1) ( ) = 2(1-1) = 2х7 - 27 dx - зму: от -2(x-1) he — 2 (1 —1) Ə d